### v2

# Hypergraph Neural Sheaf Diffusion: A Symmetric Simplicial Set Framework for Higher-Order Learning

Abstract
The absence of intrinsic adjacency relations and orientation systems in hypergraphs creates fundamental challenges for constructing sheaf Laplacians of arbitrary degrees. We resolve these limitations through symmetric simplicial sets derived directly from hypergraphs, which encode all possible oriented subrelations within each hyperedge as ordered tuples. This construction canonically defines adjacency via facet maps while inherently preserving hyperedge provenance. We establish that the normalized degree zero sheaf Laplacian on our induced symmetric simplicial set reduces exactly to the traditional graph normalized sheaf Laplacian when restricted to graphs, validating its mathematical consistency with prior graph-based sheaf theory. Furthermore, the induced structure preserves all structural information from the original hypergraph, ensuring that every multi-way relational detail is faithfully retained. Leveraging this framework, we introduce Hypergraph Neural Sheaf Diffusion (HNSD), the first principled extension of Neural Sheaf Diffusion (NSD) to hypergraphs. HNSD operates via normalized degree zero sheaf Laplacians over symmetric simplicial sets, resolving orientation ambiguity and adjacency sparsity inherent to hypergraph learning. Experimental evaluations demonstrate HNSD’s competitive performance across established benchmarks.
Index Terms:
Hypergraph Neural Networks, Hypergraph Laplacian, Symmetric Simplicial Sets, Sheaf Laplacian, Cellular Sheaf Theory
I Introduction
Hypergraphs are mathematical structures that effectively represent higher-order relationships among entities through hyperedges, each of which may connect more than two nodes simultaneously[1]. As many real-world datasets naturally form hypergraphs[2, 3, 4], constructing effective neural network architectures on hypergraphs has become crucial for learning from such complex data[5, 6, 7]. A foundational approach for relational data involves designing neural architectures around Laplacian operators, which encode relational structures into learnable diffusion processes.
Graphs—viewed as 1-dimensional hypergraphs where hyperedges connect exactly two nodes—provide a natural foundation for Laplacian-based architectures. Graphs satisfy two critical structural conditions: (1) adjacency between nodes is canonically defined through pairwise edges, and (2) no total ordering of nodes is required to define the Laplacian. This structural simplicity allows graph Laplacians to encode relational patterns effectively, forming the basis for Graph Convolutional Networks (GCNs)[8, 9], which have shown significant success, especially for homophilic graph datasets[10, 11, 12, 13]. Cellular sheaf theory [14] enhances this framework by equipping graphs with additional algebraic-topological structure, assigning vector spaces to nodes and edges, and defining consistency constraints between them. This leads to the degree 0 sheaf Laplacian [15], a generalization of the standard graph Laplacian, which underpins the development of Neural Sheaf Diffusion(NSD) models on graphs [16, 17, 18]. Unlike traditional GCNs, NSD leverages the expressive power of cellular sheaves to address two major limitations: poor performance on heterophilic graphs and the oversmoothing problem, where node representations become indistinguishable in deep architectures.
However, hypergraphs pose inherent challenges due to their structural complexity, making it extremely difficult to define sheaf Laplacians of various degrees as is done for graphs. In graphs, which model only 1-dimensional relations, the degree 0 sheaf Laplacian suffices to capture all connectivity patterns. In contrast, hypergraphs naturally encode high-dimensional relations, necessitating the definition of sheaf Laplacians of arbitrary degrees to represent their structure fully. Achieving this requires a well-defined notion of adjacency between hyperedges. Yet, by design, hypergraphs lack explicit adjacency among hyperedges-even with many hyperedges present, their multi-node nature often leads to sparse or undefined adjacency patterns. This sparsity can trivialize Laplacian operators, rendering them ineffective for learning tasks. These limitations highlight a critical gap: while Laplacian-based architectures are central to relational learning, their direct application to hypergraphs remains hindered by structural mismatches, motivating novel approaches to reconcile hypergraph complexity with effective operator design.
$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$e$$e^{\prime}$
(a) Hypergraph
$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$
(b) Induced simplicial complex
$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$
(c) Induced symmetric simplicial set
Figure 1: Comparison of different methods for representing and analyzing hypergraphs. (a) Hypergraphs naturally capture high-dimensional relationships among entities but lack explicit adjacency by design. Even with numerous hyperedges, their multi-node structure often leads to sparse or undefined adjacency, which can trivialize sheaf Laplacian of high degrees and limit their effectiveness for learning tasks. (b) Induced simplicial complex is obtained by adjoining all hyperedge subsets as simplices suffer critical ambiguities. For example, the simplex $\{v_{2},v_{3}\}$ could derive from: (1) a subrelation of $e$, (2) a subrelation of $e^{\prime}$, or (3) exist as an independent hyperedge. The induced simplicial complex treats all three cases identically, irrevocably erasing contextual origins. This loss of provenance makes hypergraph reconstruction impossible. Furthermore, sheaf Laplacians further inherit artificial dependencies on arbitrary node orderings. (c) The induced symmetric simplicial set resolves these limitations by systematically adjoining all tuples from each hyperedge while preserving hyperedge provenance through explicit tuple labeling. For example, the oriented simplex $[v_{2},v_{3}]_{e}$ maintains traceable origins as a subrelation of hyperedge $e$, while $[v_{2},v_{3}]_{e^{\prime}}$ is explicitly tagged as deriving from hyperedge $e^{\prime}$. This ensures full structural reconstruction of the original hypergraph while defining adjacency canonically via facet maps.
Given the structural challenges of hypergraphs, a natural strategy to define sheaf Laplacians involves embedding hypergraphs into simplicial complexes [19] by adjoining all subsets of each hyperedge as simplices(Fig. 1(b)). This approach interprets simplex as a possible subrelation of each hyperedge represented by a set. The induced structure inherently supports adjacency: two simplices are adjacent if they share a common facet or cofacet, creating a dense network of relationships [20, 21, 22]. Combined with the ability to induce orientations via a total node order, simplicial complexes initially appear well-suited for defining sheaf Laplacians of various degrees. However, this strategy introduces two critical limitations. First, hyperedge-specific structural information is lost: $\{v_{2},v_{3}\}$ in Fig. 1(b) become indistinguishable whether it represents possible subrelations of hyperedge $e$, hyperedge $e^{\prime}$, or independently observed hyperedge. This ambiguity erases contextual relationships between subrelations within hyperedge and the original hyperedges, making it impossible to reconstruct the original hypergraph; and (2) dependence on an arbitrary total node order to define orientations [23].
In this work, we address these limitations by introducing a symmetric simplicial set explicitly constructed from hypergraphs. Our framework collects all possible tuples in each hyperedge with its originating hyperedge(Fig. 1(c)). These tuples, which we term simplices, encode all possible oriented subrelations derived from the observed hyperedges while preserving their hyperedge-specific context. The induced structure inherently defines adjacency via facet maps and supports sheaf Laplacians of arbitrary degrees without requiring auxiliary node orderings. Crucially, our construction generalizes traditional graph-based sheaf theory: the normalized degree 0 sheaf Laplacian on the symmetric simplicial set reduces exactly to the traditional normalized sheaf Laplacian when restricted to the graph. This equivalence validates our framework as a principled extension of sheaf theory to hypergraphs. Leveraging this foundation, we develop Hypergraph Neural Sheaf Diffusion (HNSD), which implements NSD on hypergraphs via normalized degree 0 sheaf Laplacian over the induced symmetric simplicial set. Our key contributions can be summarized as follows:
  • We introduce a symmetric simplicial set construction from a hypergraph, thus enabling the formulation of high-degree sheaf Laplacians on a hypergraph.
  • We demonstrate two fundamental properties of the induced symmetric simplicial set: induced symmetric simplicial set preserves all structural information of the original hypergraph, and the normalized degree 0 sheaf Laplacian is equal to the traditional normalized sheaf Laplacian in the case of graphs.
  • We introduce new architecture HNSD, the first extension of NSD to hypergraphs, implemented via the normalized degree 0 sheaf Laplacian on the induced symmetric simplicial set. HNSD resolves fundamental orientation and adjacency ambiguities in higher-order learning while achieving competitive performance across established hypergraph benchmarks.
The remainder of this paper is structured as follows. Section II reviews related work on the sheaf Laplacians on graphs, hypergraphs, and the symmetric simplicial sets derived from hypergraphs. Section III provides preliminaries for sheaf Laplacians on symmetric simplicial sets. In Section IV, we establish a systematic framework for defining sheaf Laplacians on hypergraphs through the construction of symmetric simplicial sets directly from hypergraph structures, detailing the categorical construction of adjacency relations and orientation-preserving face maps. We then introduce the HNSD framework, implemented via normalized degree 0 sheaf Laplacians on the induced symmetric simplicial set. Finally, experimental evaluations conducted on real-world datasets are presented in Section V, and concluding remarks are offered in Section VI.
TABLE I: Summary of symbols used in the paper
Notation Description
$\mathbb{N}$ the set of non-negative integers
$[n]$$\{0,1,\cdots,n\}$ for $n\in\mathbb{N}$
$\mathfrak{S}_{n}$ all bijections from $[n]$ to $[n]$
$\|A\|$ $|A|-1$ for a set $A$
$2^{B}$ the set of all subsets of $B$
$B^{A}$ the set of all functions from $A$ to $B$
$\mathrm{Id}$ the identity map any set to itself
II Related works
II-A Sheaf learning on graphs
The graph Laplacian serves as the foundational tool for constructing neural network architectures [9, 8]. Despite their remarkable success in homophilic graph learning, GCNs face significant challenges in heterophilic graphs [24] and suffer from oversmoothing with increasing network depth [25, 26]. These limitations stem from the fundamental constraints of traditional graph Laplacian-based diffusion processes. Recent advances addressed these issues through geometric reinterpretations using cellular sheaf theory [14]. The foundational implementation emerged with Sheaf Neural Networks [27], which replace standard graph Laplacians with sheaf Laplacians to enable more sophisticated feature transport. Subsequent developments introduced learnable cellular sheaves [17, 18], demonstrating that principled diffusion processes guided by sheaf geometry can simultaneously mitigate oversmoothing and improve performance on heterophilic graphs.
II-B Hypergraph networks
In hypergraph learning, hypergraph Laplacians derived from incidence matrices have been used to construct neural networks on hypergraphs [5, 28, 29, 30, 31]. These incidence matrix-derived operators have been further extended through sheaf-theoretic frameworks, where node-level analysis is enhanced via sheaf Laplacians that model feature transport through hyperedge-mediated interactions [32]. While simplicial complexes impose additional structural constraints compared to general hypergraphs, recent work demonstrated their value in neural architectures through orientation-aware diffusion processes. Specifically, degree $n$ sheaf Laplacians have been adapted to operate on simplicial complexes [21, 22, 20], governing how features at $n$-dimensional simplices diffuse along their adjacent simplices.
II-C Symmetric Simplicial Sets from Hypergraphs
A categorical framework established a systematic relationship between ordered hypergraphs and simplicial sets through functorial construction, where ordered hypergraphs are mapped to their canonical simplicial set representations [33]. This approach was later extended to general hypergraphs via an explicit functorial construction that generates symmetric simplicial sets [34], thereby providing a concrete categorical mechanism to study hypergraph data through homological algebra.
III Preliminaries
In this section, we first review the definition of the sheaf Laplacian on graphs. We then introduce hypergraphs and their notion of isomorphism. Next, we define ordered simplicial complexes and symmetric simplicial sets, both of which organize collections of oriented $n$-dimensional simplices. Finally, we explain how adjacency between simplices is defined via facet maps, which we use to extend the sheaf Laplacian to both ordered simplicial complexes and symmetric simplicial sets.
III-A Sheaf Laplacian on graphs
A graph is a mathematical structure representing pairwise (1-dimensional) relations among entities. Formally, a graph $G=(V(G),E(G),f_{G})$ consists of a node set $V(G)$, an edge set $E(G)$ and a labeling function $f_{G}:E(G)\to 2^{V(G)}$ satisfying $|f_{G}(e)|=2$ for each edge $e\in E(G)$. A node $v$ is said to be incident to an edge $e$, denoted by $v\unlhd e$, if and only if $v\in f_{G}(e)$.
In the context of graph-based data analysis, vector-valued features are typically assigned to nodes. A central question is how these node features diffuse along edges, and this diffusion is precisely captured by the notion of a cellular sheaf and sheaf Laplacian [14, 15].
Definition 1
Given a graph $G$, a cellular sheaf $(G,\mathcal{F})$ consists of the following data:
  • For each node $v\in V(G)$ and each edge $e\in E(G)$, real vector spaces $\mathcal{F}(v)$ and $\mathcal{F}(e)$, called the stalks at node $v$ and edge $e$, respectively.
  • For each $v\unlhd e$, a linear map
    $\mathcal{F}(v\unlhd e):\mathcal{F}(v)\rightarrow\mathcal{F}(e).$
A 0-cochain $\mathbf{x}$ on the graph $G$, denoted by $\mathbf{x}=(x_{v})_{v\in V}$ is an element of the direct sum of stalks over all nodes of $G$. We denote the set of all 0-cochains as $C^{0}(G,\mathcal{F})$. For a given 0-cochain $\mathbf{x}$, the $v$-component of the sheaf Laplacian, denoted by $L_{\mathcal{F}}(\mathbf{x})_{v}$, is defined as
$L_{\mathcal{F}}(\mathbf{x})_{v}=\sum_{v,u\unlhd e}\mathcal{F}^{*}(v\unlhd e)\left(\mathcal{F}(v\unlhd e)(x_{v})-\mathcal{F}(u\unlhd e)(x_{u})\right).$ (1)
Additionally, the $v$-component of the diagonal blocks, denoted by $D_{\mathcal{F}}(\mathbf{x})_{v}$, is defined as
$\sum_{\{e\mid v\unlhd e\}}\mathcal{F}^{*}(v\unlhd e)\mathcal{F}(v\unlhd e)(x_{v}).$ (2)
Finally, the $v$-component of the normalized sheaf Laplacian, denoted by $\mathcal{L}_{\mathcal{F}}(\mathbf{x})_{v}$, is defined by
$\left((D_{\mathcal{F}})^{-\frac{1}{2}}L_{\mathcal{F}}(D_{\mathcal{F}})^{-\frac{1}{2}}\right)(\mathbf{x})_{v}.$ (3)
A cellular sheaf $\mathcal{F}$ on $G$ enables the diffusion of a node feature $x_{v}$ to an adjacent node $w$ via $\mathcal{F}^{*}(w\unlhd e)\mathcal{F}(v\unlhd e)(x_{v})$. A 0-cochain $\mathbf{x}$ represents the collection of node features $x_{v}$ for each node $v$. The operator $L_{\mathcal{F}}(\mathbf{x})$ describes how the collection of node features $\mathbf{x}$ diffuses to adjacent nodes according to the cellular sheaf $\mathcal{F}$, thereby generalizing the graph Laplacian. Similarly, $\mathcal{L}_{\mathcal{F}}$ extends the concept of the normalized graph Laplacian.
III-B Hypergraphs
A hypergraph generalizes the notion of a graph by capturing not only pairwise relationships but also higher-order, multidimensional interactions among entities.
Definition 2
A hypergraph $H=(V(H),E(H),f_{H})$ consists of the following data :
  1. 1.
    $V(H)$ is the node set of $H$
  2. 2.
    $E(H)$ is the hyperedge set of $H$
  3. 3.
    $f_{H}:E(H)\to 2^{V(H)}\backslash V(H)$, the labeling function of $H$.
For each hyperedge $e\in E(H)$, we denote its dimension by $\|e\|:=\|f_{H}(e)\|$. Two hypergraphs $H$ and $H^{\prime}$ are isomorphic [1] if there exist bijections $a:V(H)\to V(H^{\prime})$ and $b:E(H)\to E(H^{\prime})$ satisfying:
$a_{*}\circ f_{H}=f_{H^{\prime}}\circ b,(a^{-1})_{*}\circ f_{H^{\prime}}=f_{H}\circ b^{-1}$
where the induced map $a_{*}:2^{V(H)}\to 2^{V(H^{\prime})}$ is defined as
$a_{*}(\{v_{0},\cdots,v_{n}\}):=\{a(v_{0}),\cdots,a(v_{n})\}.$
A graph is a special case of a hypergraph in which every hyperedge has dimension one; equivalently, $\|e\|=1$ for all $e\in E(H)$. Two hypergraphs are isomorphic if the patterns of how entities participate together in high-order relations are the same in both, regardless of the names of the entities.
III-C Ordered simplicial complexes
A simplicial complex $X$ [35] is a collection of unoriented $n$-dimensional objects, called simplices of $X$, represented by sets. Formally, let $X\subseteq 2^{S}$ for some finite set $S$. The set $X$ is called a simplicial complex if $\tau\in X$ and $\sigma\subset\tau$, then $\sigma\in X$. An element $\sigma\in X$ is referred to as $n$-simplex if $\|\sigma\|=n$. We denote $X_{n}$ as the set of $n$-simplices of $X$. The vertex set of the simplicial complex is defined by $V(X):=\underset{\sigma\in X}{\bigcup}\sigma$. A common method for uniformly orienting the simplices of a simplicial complex is by imposing a total order on the vertex set $V(X)$.
Definition 3
Let $X$ be a simplicial complex.
  1. 1.
    $(X,<)$ is called an ordered simplicial complex if $<$ is a total order on the vertex set $V(X)$.
  2. 2.
    Suppose $(X,<)$ is an ordered simplicial complex.
    1. (a)
      The $i$th facet map, denoted by $d^{n}_{i}:X_{n}\to X_{n-1}$, is defined by
      $d^{n}_{i}(\{v_{j_{0}},\cdots,v_{j_{n}}\}):=\{v_{j_{0}},\cdots,\widehat{v_{j_{i}}},\cdots,v_{j_{n}}\}$
      where the vertices satisfy $v_{j_{0}}<\cdots<v_{j_{n}}$. If $\sigma=d_{i}^{n}(\tau)$ for some $\tau$, we say $\sigma$ is a facet of $\tau$ (or equivalently, $\tau$ is a cofacet of $\sigma$), denoted by $\sigma\unlhd\tau$.
    2. (b)
      For simplices satisfying $\sigma\unlhd\tau$, the signed incidence of $\sigma,\tau$ [14], denoted by $[\sigma:\tau]$, is defined by $(-1)^{i}$ where $\sigma=d^{n}_{i}(\tau)$.
When $\sigma\unlhd\tau$, the signed incidence $[\sigma:\tau]$ indicates whether the orientation of the facet $\sigma$ is consistent, denoted by $+1$, or reversed, denoted by $-1$, relative to the orientation of $\tau$.
Example 1
Let $V$ be a finite set and consider the simplicial complex $X=2^{V}\setminus\emptyset$ with vertex set $V(X)=V$. An illustration of this simplicial complex $X$ for $V=\{v_{0},v_{1},v_{2}\}$ is provided in Fig. 2. Suppose that $<$ is a total order on $V(X)$ defined by $v_{0}<v_{1}<v_{2}$. Then we have the following facet maps and signed incidences:
  • $d^{2}_{0}(\{v_{0},v_{1},v_{2}\})=\{v_{1},v_{2}\}$ and
    [$[\{v_{1},v_{2}\}:\{v_{0},v_{1},v_{2}\}]=+1.$
  • $d^{2}_{1}(\{v_{0},v_{1},v_{2}\})=\{v_{0},v_{2}\}$ and
    [$[\{v_{0},v_{2}\}:\{v_{0},v_{1},v_{2}\}]=-1.$
  • $d^{2}_{2}(\{v_{0},v_{1},v_{2}\})=\{v_{0},v_{1}\}$ and
    [$[\{v_{0},v_{1}\}:\{v_{0},v_{1},v_{2}\}]=+1.$
[…]
Figure 2: Illustration of the simplicial complex $2^{V}\setminus\emptyset$ for the vertex set $V=\{v_{0},v_{1},v_{2}\}$ with the total order $v_{0}<v_{1}<v_{2}$. The 1-simplices $\{v_{0},v_{1}\}$, $\{v_{0},v_{2}\}$, and $\{v_{1},v_{2}\}$ are represented by the sides of the 2-simplex $\{v_{0},v_{1},v_{2}\}$, which is depicted as the gray triangle. Each side is a facet of the gray triangle, and their corresponding signed incidences are indicated accordingly.
III-D Symmetric simplicial sets
A symmetric simplicial set $X$ [36] is a collection of oriented $n$-dimensional objects, called simplices of $X$, represented by tuples, such that each simplex has well-defined facets.
Definition 4
Let $X=\{X_{n}\}_{n\in\mathbb{N}}$ be a collection of finite sets.
  1. 1.
    The collection $X$ is called a symmetric simplicial set if it is equipped with a family of functions
    $\{X(\mu):X_{n}\to X_{m}\}_{\{\mu:[m]\to[n]\}}$
    satisfying if $\mu:[m]\to[n],\nu:[n]\to[p]$ then
    $X(\nu\circ\mu)=X(\mu)\circ X(\nu).$ (4)
    An element of $X_{n}$ is called $n$-simplex of $X$. We say two $n$-simplices $\sigma,\sigma^{\prime}$ are equivalent if
    $X(g)(\sigma^{\prime})=\sigma$ (5)
    for some $g\in\mathfrak{S}_{n}$. We denote the equivalence class of $\sigma$ by $[\sigma]$.
  2. 2.
    For an $n$-simplex $\sigma$ of $X$, $i\in[n]$,
    1. (a)
      the $i$th vertex of $\sigma$, denoted by $\sigma_{i}$, is defined as
      $X\left((i)_{[n]}\right)(\sigma)$
      where $(i)_{[n]}:[0]\to[n]$ is a function with $(i)_{[n]}(0)=i$. $(\sigma_{0},\cdots,\sigma_{n})$ is called the tuple representation of $\sigma$.
    2. (b)
      the $i$th facet of $\sigma$, denoted by $d^{n}_{i}(\sigma)$, is defined as
      $X\left(\delta^{n}_{i}\right)(\sigma)$
      where $\delta^{n}_{i}:[n-1]\to[n]$ is the unique order preserving injection from $\{0,\cdots,n-1\}$ to $\{0,\cdots,\hat{i},\cdots,n\}$. If $\sigma=d_{i}^{n}(\tau)$ for some $\tau$, we say $\sigma$ is a facet of $\tau$ (or equivalently, $\tau$ is a cofacet of $\sigma$), denoted by $\sigma\unlhd\tau$.
    3. (c)
      For simplices $\sigma\unlhd\tau$, the signed incidence of $\sigma,\tau$, denoted by $[\sigma:\tau]$, is defined by $(-1)^{i}$ for $\sigma=d^{n}_{i}(\tau)$.
    4. (d)
      Let $\sigma^{n}_{i}:[n+1]\to[n]$ be the surjection defined by
      $\sigma^{n}_{i}:=\begin{cases}j,&\text{if $j\leq i$}\\ j-1,&\text{if $j>i$.}\end{cases}$
    An $n$-simplex $\sigma$ is said to be degenerate if it lies in the image of $X(\sigma_{i}^{n})$ for some $i\in[n]$. Otherwise, $\sigma$ is said to be nondegenerate.
An $n$-simplex $\sigma$ has a tuple representation $(\sigma_{0},\cdots,\sigma_{n})$. Two $n$-simplices $\sigma,\sigma^{\prime}$ are equivalent if $(\sigma_{0},\cdots,\sigma_{n})$ is equal to $(\sigma^{\prime}_{g(0)},\cdots,\sigma^{\prime}_{g(n)})$ for some permutation $g\in\mathfrak{S}_{n}$. The equivalence class $[\sigma]$ is represented by a set $\{\sigma_{0},\cdots,\sigma_{n}\}$. The facet $d_{i}^{n}(\sigma)$ is a $(n-1)$ simplex with tuple representation $(\sigma_{0},\cdots,\widehat{\sigma_{i}},\cdots,\sigma_{n})$. The nondegeneracy of $\sigma$ implies $\sigma_{i}\neq\sigma_{j}$ for all $i\neq j$. These combinatorial structures of $\sigma$ are systematically encoded via maps $\mu:[m]\to[n]$, without explicitly referencing the tuple representation of $\sigma$.
Example 2
Let $V$ be a finite set. A $V$-simplex $\Delta[V]$ [34] is a symmetric simplicial set defined by the following data:
  • The simplices are given by the family
    $\Delta[V]=\{\Delta[V]_{n}:=V^{[n]}\}_{n\in\mathbb{N}}.$
  • For any function $\mu:[m]\to[n]$, the induced map $\Delta[V](\mu):\Delta[V]_{n}\to\Delta[V]_{m}$ is defined as
    $\Delta[V](\mu)((v_{i_{0}},\cdots,v_{i_{n}})_{V}):=(v_{i_{\mu(0)}},\cdots,v_{i_{\mu(m)}})_{V}.$ (6)
An illustrative example of this symmetric simplicial set structure is provided in Fig. 3. It can be verified that $\Delta[V]$ satisfies the compatibility condition (Equation (4)), thus confirming that $\Delta[V]$ is indeed a symmetric simplicial set.
  • Tuple representation of $(v_{j_{0}},\cdots,v_{j_{n}})_{V}$ is $(v_{j_{0}},\cdots,v_{j_{n}})$.
  • The $i$th vertex of the simplex $(v_{j_{0}},\cdots,v_{j_{n}})_{V}$ is $v_{j_{i}}$.
  • The $i$th facet is obtained by
    $d^{n}_{i}\left((v_{j_{0}},\cdots,v_{j_{n}})_{V}\right)=(v_{j_{0}},\cdots,\widehat{v_{j_{i}}}\cdots,v_{j_{n}})_{V}.$
  • A simplex $(v_{j_{0}},\cdots,v_{j_{n}})_{V}$ is degenerate if and only if $v_{j_{i}}=v_{j_{i^{\prime}}}$ for some $i\neq i^{\prime}$. In particular, the degenerate $n$-simplex $(v,\cdots,v)_{V}$ is called the $v$ of multiplicity $(n+1)$.
Hence, an $n$-simplex in $\Delta[V]$ can be naturally interpreted as an ordered $(n+1)$-tuple of elements in $V$.
In Examples 1 and 2, we constructed an ordered simplicial complex and a symmetric simplicial set, respectively, from a given finite vertex set $V$. These two constructions differ as follows:
  • A simplex in $2^{V}\setminus\emptyset$ is represented as an unordered subset of $V$, whereas a simplex in $\Delta[V]$ is represented as an ordered tuple in $V$.
  • $2^{V}\setminus\emptyset$ requires a total order on the vertex set $V$ to define the $i$th facet of a simplex explicitly. In contrast, $\Delta[V]$ does not depend on any choice of total order on $V$.
  • In $2^{V}\setminus\emptyset$, no $n$-simplices exist for $n>|V|-1$. In contrast, $\Delta[V]$ contains $n$-simplices for every $n\in\mathbb{N}$, due to the allowance of degenerate simplices.
[…]
Figure 3: Illustration of the symmetric simplicial set $\Delta[V]$ for $V=\{v_{0},v_{1},v_{2}\}$ described by 2-dimensional nondegenerate simplex $(v_{0},v_{1},v_{2})_{V}$. The 1-simplices $(v_{0},v_{1})_{V},(v_{0},v_{2})_{V}$ and $(v_{1},v_{2})_{V}$ are depicted by directed edges, indicating the ordered 2-tuple structures. The 2-simplex $(v_{0},v_{1},v_{2})_{V}$ is represented by the gray triangle oriented in a counterclockwise direction. Each directed edge is a facet of the gray triangle, and their corresponding signed incidences are indicated accordingly.
III-E Sheaf Laplacians
In this subsection, we assume $X$ is either an ordered simplicial complex or a symmetric simplicial set. We define upper and lower adjacency relations between simplices of the same dimension in $X$ motivated by  [37].
Definition 5
Let $\sigma,\sigma^{\prime}\in X_{n}$ for some $n\in\mathbb{N}$.
  1. 1.
    The simplices $\sigma$ and $\sigma^{\prime}$ are said to be upper adjacent if there exists a common cofacet $\tau\in X_{n+1}$ such that $\sigma\unlhd\tau$ and $\sigma^{\prime}\unlhd\tau$.
  2. 2.
    The simplices $\sigma$ and $\sigma^{\prime}$ are said to be lower adjacent if there exists a common facet $\mu\in X_{n-1}$ such that $\mu\unlhd\sigma$ and $\mu\unlhd\sigma^{\prime}$.
In other words, two simplices are upper adjacent if they share a common cofacet, and they are lower adjacent if they share a common facet.
Example 3
Consider $\Delta[V]$ for $V=\{v_{0},v_{1},v_{2}\}$. Fig. 4 illustrates upper and lower adjacency relations between simplices:
  • The simplices $(v_{0})_{V}$ and $(v_{1})_{V}$ are upper adjacent since both are facets of the common simplex $(v_{0},v_{1})_{V}$.
  • The simplices $(v_{0},v_{1})_{V}$ and $(v_{1},v_{2})_{V}$ are lower adjacent since both share the common facet $(v_{1})_{V}$.
[…]
Figure 4: Illustration of adjacency relations in the symmetric simplicial set $\Delta[V]$ for $V=\{v_{0},v_{1},v_{2}\}$. The simplices $(v_{0})_{V}$ and $(v_{1})_{V}$ are upper adjacent, sharing the common cofacet $(v_{0},v_{1})_{V}$. Similarly, the simplices $(v_{0},v_{1})_{V}$ and $(v_{1},v_{2})_{V}$ are lower adjacent, sharing the common facet $(v_{1})_{V}$.
To utilize simplices for data analysis, it is essential to assign vector-valued features to each simplex. Moreover, given rules that transfer features from an $n$-simplex to its facets and vice versa, it becomes possible to diffuse features between adjacent simplices. Such feature diffusion mechanisms are formally captured by the concept of a cellular sheaf [34].
Definition 6
A cellular sheaf $(X,\mathcal{F})$ of degree $m$ consists of the following data:
  • For $n\in[m]$, $n$-simplex $\sigma\in X_{n}$, a $\mathbb{R}$-vector space $\mathcal{F}(\sigma)$, called the stalk at $\sigma$.
  • For $n\in[m]$, $n$-simplex $\sigma\in X_{n}$ and facet $d_{i}^{n}(\sigma)$, a linear map
    $\mathcal{F}(d_{i}^{n}(\sigma)\unlhd\sigma):\mathcal{F}(d_{i}^{n}(\sigma))\rightarrow\mathcal{F}(\sigma)$
    satisfying the following compatibility conditions:
    $\mathcal{F}\left(d_{j}^{n-1}(d_{i}^{n}(\sigma))\unlhd d_{i}^{n}(\sigma)\right)\circ\mathcal{F}\left(d_{i}^{n}(\sigma)\unlhd\sigma\right)$
    $=\mathcal{F}\left(d_{i-1}^{n-1}(d_{j}^{n}(\sigma))\unlhd d_{j}^{n}(\sigma)\right)\circ\mathcal{F}\left(d_{j}^{n}(\sigma)\unlhd\sigma\right)$ (7)
    for every $n\in[m]$, $i\in[n]$, $j<i$, and $\sigma\in X_{n}$.
Elements of $\mathcal{F}(\sigma)$ are called features at simplex $\sigma$.
A cellular sheaf $\mathcal{F}$ of degree $m$ on $X$ assigns a vector space to each $n$-simplex for $0\leq n\leq m$, together with linear maps between these vector spaces corresponding to the facet relations. When $m$ is sufficiently large, we may omit specifying its value. Intuitively, the map $\mathcal{F}^{*}(d_{i}^{n}(\sigma)\unlhd\sigma)$ describes how features at the simplex $\sigma$ are restricted to its $i$th facet, while the map $\mathcal{F}(d_{i}^{n}(\sigma)\unlhd\sigma)$ describes how features from the $i$th facet are propagated upward to $\sigma$.
Using the cellular sheaf structure, a feature $x_{\sigma}\in\mathcal{F}(\sigma)$ at simplex $\sigma$ can be diffused to an adjacent simplex $\sigma^{\prime}$ as follows. Suppose $\sigma$ and $\sigma^{\prime}$ are upper adjacent with common cofacet $\tau$. Then, the feature $x_{\sigma}$ can be diffused to $\sigma^{\prime}$ by:
$(-1)^{[\sigma:\tau]+[\sigma^{\prime}:\tau]}\mathcal{F}^{*}(\sigma^{\prime}\unlhd\tau)\,\mathcal{F}(\sigma\unlhd\tau)(x_{\sigma})\in\mathcal{F}(\sigma^{\prime}).$
Similarly, if $\sigma$ and $\sigma^{\prime}$ are lower adjacent with common facet $\mu$, the diffusion rule becomes:
$(-1)^{[\mu:\sigma]+[\mu:\sigma^{\prime}]}\mathcal{F}(\mu\unlhd\sigma^{\prime})\,\mathcal{F}^{*}(\mu\unlhd\sigma)(x_{\sigma})\in\mathcal{F}(\sigma^{\prime}).$
These feature diffusions between adjacent simplices are illustrated in Fig. 5. Aggregating all such possible feature diffusions according to these rules results in the construction of the sheaf Laplacian.
${\mathcal{F}(\tau)}$${\mathcal{F}(\sigma)}$${\mathcal{F}(\sigma^{\prime})}$${\mathcal{F}(\mu)}$$\scriptstyle{(-1)^{[\sigma^{\prime}:\tau]}\mathcal{F}^{*}(\sigma^{\prime}\unlhd\tau)}$$\scriptstyle{(-1)^{[\sigma:\tau]}\mathcal{F}(\sigma\unlhd\tau)}$$\scriptstyle{(-1)^{[\mu:\sigma]}\mathcal{F}^{*}(\mu\unlhd\sigma)}$$\scriptstyle{(-1)^{[\mu:\sigma^{\prime}]}\mathcal{F}(\mu\unlhd\sigma^{\prime})}$
Figure 5: Illustration of feature diffusion between adjacent simplices $\sigma$ and $\sigma^{\prime}$. If $\sigma$ and $\sigma^{\prime}$ share a common cofacet $\tau$, the feature at $\sigma$ is diffused to $\sigma^{\prime}$ via the map $(-1)^{[\sigma^{\prime}:\tau]+[\sigma:\tau]}\mathcal{F}^{*}(\sigma^{\prime}\unlhd\tau)\,\mathcal{F}(\sigma\unlhd\tau).$ Alternatively, if $\sigma$ and $\sigma^{\prime}$ share a common facet $\mu$, the feature at $\sigma$ is diffused to $\sigma^{\prime}$ via the map $(-1)^{[\mu:\sigma^{\prime}]+[\mu:\sigma]}\mathcal{F}(\mu\unlhd\sigma^{\prime})\,\mathcal{F}^{*}(\mu\unlhd\sigma).$
Definition 7
Let $\mathcal{F}$ be a cellular sheaf on $X$ and $k\in\mathbb{N}$.
  1. 1.
    A $k$-cochain, denoted by $\mathbf{x}=(x_{\sigma})_{\sigma\in X_{k}}$, is an element of the direct sum of stalks over all $k$-simplices in $X$. We call $x_{\sigma}$ the $\sigma$-component of the cochain $\mathbf{x}$. The set of all $k$-cochains is denoted by $C^{k}(X,\mathcal{F})$.
  2. 2.
    For a $k$-cochain $\mathbf{x}$, the $\sigma$-component of the degree $k$ sheaf Laplacian $L_{\mathcal{F}}^{k}(\mathbf{x})_{\sigma}$ is defined by
    $\sum_{\sigma^{\prime\prime},\tau}(-1)^{[\sigma:\tau]+[\sigma^{\prime}:\tau]}\mathcal{F}^{*}(\sigma\unlhd\tau)\mathcal{F}(\sigma^{\prime}\unlhd\tau)(x_{\sigma^{\prime}})$ (8)
    $+\sum_{\sigma^{\prime\prime},\mu}(-1)^{[\mu:\sigma]+[\mu:\sigma^{\prime\prime}]}\mathcal{F}(\mu\unlhd\sigma)\mathcal{F}^{*}(\mu\unlhd\sigma^{\prime\prime})(x_{\sigma^{\prime\prime}})$
    where
    • $\sigma^{\prime}$ ranges over all simplices upper adjacent to $\sigma$
    • $\sigma^{\prime\prime}$ ranges over all simplices lower adjacent to $\sigma$
    • $\tau$ ranges over all common cofacets of $\sigma$ and $\sigma^{\prime}$
    • $\mu$ ranges over all common facets of $\sigma$ and $\sigma^{\prime\prime}$.
    Additionally, the $\sigma$-component of the diagonal blocks, denoted by $D^{k}_{\mathcal{F}}(\mathbf{x})_{\sigma}$, is defined as
    $\sum_{\{\tau\mid\sigma\unlhd\tau\}}\mathcal{F}^{*}(\sigma\unlhd\tau)\mathcal{F}(\sigma\unlhd\tau)(x_{\sigma})$ (9)
    +$+\sum_{\{\mu\mid\mu\unlhd\sigma\}}\mathcal{F}(\mu\unlhd\sigma)\mathcal{F}^{*}(\mu\unlhd\sigma)(x_{\sigma}).$
  3. 3.
    For a $k$-cochain $\mathbf{x}$, the $\sigma$-component of the normalized degree $k$ sheaf Laplacian, denoted by $\mathcal{L}^{k}_{\mathcal{F}}(\mathbf{x})_{\sigma}$, is defined by
    $\left((D^{k}_{\mathcal{F}})^{-\frac{1}{2}}L^{k}_{\mathcal{F}}(D^{k}_{\mathcal{F}})^{-\frac{1}{2}}\right)(\mathbf{x})_{\sigma}.$ (10)
If $X$ is an ordered simplicial complex with a specified total order $<$, we explicitly denote the sheaf Laplacian as $L_{\mathcal{F},<}^{k}$.
Given a cellular sheaf $\mathcal{F}$, a $k$-cochain provides an assignment of features to every $k$-simplex in $X$. The degree $k$ sheaf Laplacian characterizes the diffusion of features from each $k$-simplex to its adjacent simplices through $\mathcal{F}$.
Example 4
Let $X=\Delta[V]$ be a $V$-simplex and $\mathcal{F}$ be a cellular sheaf on $X$. Suppose $\mathbf{x}$ is a 0-cochain and $v,w\in V$. Then the $(v)_{V}$-component of the degree 0 sheaf Laplacian can be computed as follows.
  • The simplices $(v)_{V}$ and $(w)_{V}$ are upper adjacent with common cofacets $(v,w)_{V}$ and $(w,v)_{V}$, satisfying the signed incidences:
    1. 1.
      $[(v)_{V}:(v,w)_{V}]=[(w)_{V}:(w,v)_{V}]=-1$
    2. 2.
      $[(w)_{V}:(v,w)_{V}]=[(v)_{V}:(w,v)_{V}]=+1$.
  • The simplex $(v)_{V}$ is also upper adjacent to itself through the common cofacets $(v,w)_{V}$ and $(w,v)_{V}$, satisfying the signed incidences:
    1. 1.
      $[(v)_{V}:(v,w)_{V}]=-1$
    2. 2.
      $[(v)_{V}:(w,v)_{V}]=+1$.
Hence, using Equation (8), the $(v)_{V}$-component of the degree 0 sheaf Laplacian is expressed as:
$\sum_{w\in V}\mathcal{F}^{*}((v)_{V}\unlhd(v,w)_{V})\mathcal{F}((v)_{V}\unlhd(v,w)_{V})(x_{v})$
$+$ $\sum_{w\in V}\mathcal{F}^{*}((v)_{V}\unlhd(w,v)_{V})\mathcal{F}((v)_{V}\unlhd(w,v)_{V})(x_{v})$
-- $\sum_{w\in V}\mathcal{F}^{*}((v)_{V}\unlhd(v,w)_{V})\mathcal{F}((w)_{V}\unlhd(v,w)_{V})(x_{w})$
-- $\sum_{w\in V}\mathcal{F}^{*}((v)_{V}\unlhd(w,v)_{V})\mathcal{F}((w)_{V}\unlhd(w,v)_{V})(x_{w}).$
IV Methodology
In this section, we first demonstrate that adjacency between hyperedges in general hypergraphs is not inherently well-defined. We discuss the most natural approach to resolving this issue—using the induced simplicial complex—which provides a clear notion of adjacency. However, we will identify two significant problems arising from this approach.
IV-A Problem definition
In a graph, two nodes are considered adjacent if an edge directly connects them, and two edges are adjacent if they share a common node. The degree 0 sheaf Laplacian effectively describes how node features diffuse through edges, while the degree 1 sheaf Laplacian captures the diffusion of edge features across adjacent edges. These concepts accurately reflect feature diffusion dynamics in graphs, provided there is an adequate number of edges.
However, such conditions do not directly translate to hypergraphs. In hypergraphs, adjacency between two nodes requires the existence of a dimension 1 hyperedge explicitly connecting them. Consequently, in the absence of dimension 1 hyperedges, the degree 0 sheaf Laplacian becomes zero. Similarly, two hyperedges of dimension $n$ are adjacent if there exists a hyperedge of dimension $n+1$ containing both or if one hyperedge of dimension $n-1$ is contained within the other. Crucially, having many hyperedges in total does not necessarily imply the presence of numerous adjacent hyperedges. The following example clearly illustrates this limitation.
Example 5
Consider a hypergraph $H$ defined by:
  • $V(H):=\{v_{0},v_{1},v_{2},v_{3}\}$, $E(H):=\{e_{0},e_{1},e_{2}\}$
  • $f_{H}(e_{0}):=\{v_{0},v_{1},v_{2}\}$
  • $f_{H}(e_{1}):=\{v_{0},v_{2},v_{3}\}$
  • $f_{H}(e_{2}):=\{v_{0},v_{1},v_{3}\}$.
This hypergraph is illustrated in Fig. 6. All hyperedges have dimension 2, and there are no dimension 1 hyperedges connecting any two nodes. Thus, the degree 0 sheaf Laplacian is zero. Furthermore, since no hyperedges contain others or share common subsets, no hyperedges are adjacent, leading the degree $n$ sheaf Laplacian to vanish for all $n$.
$v_{1}$$v_{0}$$v_{2}$$v_{3}$$e_{0}$$e_{1}$$e_{2}$
Figure 6: Illustration of the hypergraph described in Example 5, which has no adjacent nodes or adjacent hyperedges. This example demonstrates that having multiple hyperedges does not necessarily guarantee nonzero sheaf Laplacians. Consequently, the sheaf Laplacian in this case does not convey meaningful structural information about the hypergraph.
To address the above-mentioned problem with hypergraphs, one natural approach involves systematically inserting all subsets of intersections among hyperedges into the set of hyperedges. This construction yields an induced simplicial complex, which we denote by
$S(H):=\{S\mid S\subseteq f_{H}(e)\text{ for some }e\in E(H)\}\cup V(H).$
The resulting complex $S(H)$ thus incorporates all possible unoriented sub-relations induced from the original hyperedge structure of $H$. The vertex set $V(S(H))$ coincides with $V(H)$. By imposing a total order on $V(H)$, the sheaf Laplacian on this complex becomes well-defined, effectively resolving the adjacency issues described earlier. However, adopting this solution introduces two new significant problems, which we outline next.
  1. 1.
    First, the original hypergraph $H$ cannot necessarily be recovered from the induced simplicial complex $S(H)$. In other words, non-isomorphic hypergraphs may induce identical simplicial complexes. For example, consider two hypergraphs $H=(\{v_{0},v_{1},v_{2}\},\{e\},f_{H})$, $H^{\prime}=(\{v_{0},v_{1},v_{2}\},\{e,e^{\prime}\},f_{H^{\prime}})$ with labeling functions defined by
    $f_{H}(e)=f_{H^{\prime}}(e)=\{v_{0},v_{1},v_{2}\},\quad f_{H^{\prime}}(e^{\prime})=\{v_{0},v_{1}\}.$
    Clearly, the hypergraphs $H$ and $H^{\prime}$ are non-isomorphic. Nevertheless, their induced simplicial complexes coincide:
    $S(H)=S(H^{\prime})=2^{\{v_{0},v_{1},v_{2}\}}\setminus\emptyset.$
    This ambiguity arises because it becomes impossible to determine whether $\{v_{0},v_{1}\}$ originates as a subrelation of hyperedge $e$, derives from hyperedge $e^{\prime}$, or exists as an independent hyperedge itself.
  2. 2.
    Second, the resulting sheaf Laplacian depends explicitly on the choice of the total order assigned to the vertex set [23]. To illustrate, consider a hypergraph $H=(\{v_{0},v_{1},v_{2}\},\{e\},f_{H})$ with a single hyperedge defined as $f_{H}(e)=\{v_{0},v_{1},v_{2}\}$. Then its induced simplicial complex is:
    $S(H)=2^{\{v_{0},v_{1},v_{2}\}}\setminus\emptyset.$
    Let $<$ and $\prec$ be two distinct total orders on the vertex set $V(S(H))=\{v_{0},v_{1},v_{2}\}$, defined by $v_{0}<v_{1}<v_{2}$ and $v_{1}\prec v_{0}\prec v_{2}$. Suppose $\mathcal{F}$ is a cellular sheaf on $S(H)$ satisfying:
    ($\mathcal{F}(\{v_{1}\}\unlhd\{v_{1},v_{2}\})=\mathcal{F}(\{v_{2}\}\unlhd\{v_{1},v_{2}\})=0.$
    For simplicity, denote
    $A(\{v_{i},v_{j}\}):=\mathcal{F}(\{v_{i},v_{j}\}\unlhd\{v_{0},v_{1},v_{2}\})$
    for $i,j\in\{0,1,2\}$. Then, for a given $\mathbf{x}\in\mathcal{F}(\{v_{1},v_{2}\})$, we have:
    $L^{1}_{\mathcal{F},<}(\mathbf{x})_{\{v_{0},v_{1}\}}=A(\{v_{0},v_{1}\})^{*}A(\{v_{1},v_{2}\})(\mathbf{x}),$
    whereas under the alternative order,
    $L^{1}_{\mathcal{F},\prec}(\mathbf{x})_{\{v_{0},v_{1}\}}$ $=-A(\{v_{0},v_{1}\})^{*}A(\{v_{1},v_{2}\})(\mathbf{x})$
    $\neq L^{1}_{\mathcal{F},<}(\mathbf{x})_{\{v_{0},v_{1}\}}.$
    Thus, changing the total order of nodes leads to distinct sheaf Laplacians, specifically:
    $L^{1}_{\mathcal{F},<}\neq L^{1}_{\mathcal{F},\prec}.$
The sheaf Laplacian is well-defined on symmetric simplicial sets without requiring a total order. Thus, if we can construct a symmetric simplicial set from a given hypergraph and demonstrate that this construction preserves the hypergraph’s structural information, we can effectively resolve the aforementioned problems inherent to hypergraphs.
IV-B Symmetric simplicial set from hypergraph
Given a hyperedge $e$, the $f_{H}(e)$-simplex $\Delta[f_{H}(e)]$ is the collection of all possible oriented relations within $e$, represented by tuples, that can be predicted from the observed relation $e$. In general hypergraphs with multiple hyperedges, the $f_{H}(e)$-simplices from different hyperedge $e$ must be glued together along shared nodes. Crucially, since the original hypergraph $H$ cannot be reconstructed from $S(H)$ due to the lack of information about which hyperedge each possible relation originated from, the elements of $\Delta[f_{H}(e)]$ and $\Delta[f_{H}(e^{\prime})]$ for different hyperedges $e,e^{\prime}$ should not be identified.
For example, in Fig. 7, $(v_{2},v_{3})_{f_{H}(e)}\in\Delta[f_{H}(e)]_{1}$ and $(v_{2},v_{3})_{f_{H}(e^{\prime})}\in\Delta[f_{H}(e^{\prime})]_{1}$ share the same image $\{v_{2},v_{3}\}$ but must be treated as distinct objects. This distinction indicates $(v_{2},v_{3})_{f_{H}(e)}$ originates from $e$, while $(v_{2},v_{3})_{f_{H}(e^{\prime})}$ originates from $e^{\prime}$, hence can avoid the problem occurred in $S(H)$.
However, not all ordered relations are treated as distinct. For instance, $(v_{2},v_{2})_{f_{H}(e)}\in\Delta[f_{H}(e)]_{1}$, $(v_{2},v_{2})_{f_{H}(e^{\prime})}\in\Delta[f_{H}(e^{\prime})]_{1}$, and $(v_{2},v_{2})_{\{v_{2}\}}\in\Delta[\{v_{2}\}]$ all represent the shared node $v_{2}$ as 2-tuples. Since we glue $\Delta[f_{H}(e)],\Delta[f_{H}(e^{\prime})]$ along the shared nodes $\{v_{2},v_{3}\}$, these elements should be identified as equivalent. Extending this principle to all hyperedges, nodes, and $n$-dimensional cases yields the induced symmetric simplicial set $\Delta(H)$.
$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$e$$e^{\prime}$$(v_{0},v_{1},v_{2},v_{3})_{f_{H}(e)}$$(v_{2},v_{3},v_{4})_{f_{H}(e^{\prime})}$
Figure 7: Motivation for an induced symmetric simplicial set from hypergraph. Consider $\Delta[f_{H}(e)],\Delta[f_{H}(e^{\prime})]$ for each hyperedge $e,e^{\prime}$ and glue them along shared nodes. Figure explains $(v_{2},v_{3})_{f_{H}(e)}\in\Delta[f_{H}(e)]_{1},(v_{2},v_{3})_{f_{H}(e^{\prime})}\in\Delta[f_{H}(e^{\prime})]_{1}$ have same image $\{v_{2},v_{3}\}$, but they are embedded in $\Delta(H)_{1}$ as different 1-simplices. Since $(v_{2})_{f_{H}(e)}\in\Delta[f_{H}(e)]_{0},(v_{2})_{f_{H}(e^{\prime})}\in\Delta[f_{H}(e^{\prime})]_{0},(v_{2})_{v_{2}}\in\Delta[\{v_{2}\}]_{v_{2}}$ represent the shared node $v_{2}$, they should be considered as equal.
Definition 8
Given a hypergraph $H$, we define its induced symmetric simplicial set, denoted by $\Delta(H)=\{\Delta(H)_{n}\}_{n\in\mathbb{N}}$, as follows:
$\Delta(H)_{n}:=\left(\coprod_{e\in E(H)}\Delta[f_{H}(e)]_{n}\right)\coprod\left(\coprod_{v\in V(H)}\Delta[\{v\}]_{n}\right)\Big{/}\sim$
where $\sim$ is the equivalence relation generated by
$\Delta(H)_{n}\ni(v,\cdots,v)_{f_{H}(e)}\sim(v,\cdots,v)_{\{v\}}\in\Delta(H)_{n}$
for any $e\in E(H)$ such that $v\in f_{H}(e)$. We denote the equivalence class of $(v_{0},\cdots,v_{n})_{f_{H}(x)}$ by $[v_{0},\cdots,v_{n}]_{x}$ for $x\in V(H)\coprod E(H)$.
For $\mu:[m]\to[n]$, $\Delta(H)(\mu):\Delta(H)_{n}\to\Delta(H)_{m}$ is defined by
$\Delta(H)(\mu)([v_{i_{0}},\cdots,v_{i_{n}}]_{e}):=[v_{i_{\mu(0)}},\cdots,v_{i_{\mu(m)}}]_{e}.$ (11)
Since Equation (11) satisfies Equation (4), the resulting $\Delta(H)$ is indeed a symmetric simplicial set. We say that the simplex $[v_{i_{0}},\dots,v_{i_{\|e\|}}]_{e}\in\Delta(H)_{\|e\|}$ is a maximal nondegenerate $\|e\|$-simplex if all vertices $v_{i_{l}}$ are distinct for $l\neq l^{\prime}$.
$\Delta[f_{H}(e)]$ is embedded into $\Delta(H)$ for each hyperedge $e$. For example, $m$-simplex $(v_{j_{0}},\ldots,v_{j_{m}})_{f_{H}(e)}$ in $\Delta[f_{H}(e)]_{m}$ is embedded into $m$-simplex $[v_{j_{0}},\ldots,v_{j_{m}}]_{e}$ in $\Delta(H)$. In particular, when $v\in f_{H}(e)$, the $v$ of multiplicity $(n+1)$, $(v,\cdots,v)_{V}$, is embedded into $[v,\ldots,v]_{e}$. We identify all $v$ of multiplicity $(n+1)$ from different hyperedges as $[v,\cdots,v]_{v}$.
Example 6
Consider a hypergraph $H=(V(H),E(H),f_{H})$
  • $V(H)=\{v_{0},v_{1},v_{2},v_{3},v_{4}\},E(H)=\{e,e^{\prime}\}$
  • $f_{H}(e)=\{v_{0},v_{1},v_{2},v_{3}\},f_{H}(e^{\prime})=\{v_{2},v_{3},v_{4}\}$.
$\Delta[f_{H}(e)],\Delta[f_{H}(e^{\prime})]$ are embedded in $\Delta(H)$ glued along $v_{2},v_{3}$ (Fig.8(a)). This means that the $v_{2}$ (or $v_{3}$) of multiplicity $(n+1)$ in $\Delta[f_{H}(e)],\Delta[f_{H}(e^{\prime})]$ are embedded into the same $n$-simplex $[v_{2},\cdots,v_{2}]_{v_{2}}$ (or $[v_{3},\cdots,v_{3}]_{v_{3}}$). The special case is when $n=0$. If $v\in f_{H}(e)\cap f_{H}(e^{\prime})$, $(v)_{f_{H}(e)},(v)_{f_{H}(e^{\prime})}$ are embedded in to the same 0-simplex $[v]_{v}$. Hence
$\Delta(H)_{0}=\{[v_{0}]_{v_{0}},[v_{1}]_{v_{1}},[v_{2}]_{v_{2}},[v_{3}]_{v_{3}},[v_{4}]_{v_{4}}\}\cong V(H).$
The same arguments show that $\Delta(H)_{0}\cong V(H)$ for any hypergraph $H$.
Given the hypergraph $H$ in Example 6, we construct a hypergraph $(V,E,f)$ (Fig. 8(b)) using solely the data from $\Delta(H)$, through the following systematic procedure:
  • Node set $V$: Each 00-simplex in $\Delta(H)$ becomes a node, hence
    $V:=V(H)\cong\Delta(H)_{0}.$
  • Hyperedge set $E$: Equivalence class of maximal nondegenerate simplex (Equation (5)) becomes a hyperedge, hence
    $E=\{[[v_{0},v_{1},v_{2},v_{3}]_{e}],\ [[v_{2},v_{3},v_{4}]_{e^{\prime}}]\}\cong E(H).$
  • Labeling function $f$: $f$ sends
    1. 1.
      $[[v_{0},v_{1},v_{2},v_{3}]_{e}]\mapsto\{v_{0},v_{1},v_{2},v_{3}\}$
    2. 2.
      $[[v_{2},v_{3},v_{4}]_{e^{\prime}}]\mapsto\{v_{2},v_{3},v_{4}\}$.
We can easily demonstrate the isomorphism between hypergraphs $(V,E,f)$ and $H$, as each equivalence class of maximal nondegenerate simplices bijectively corresponds to a hyperedge in $H$, and vice versa. This reconstruction method extends naturally to arbitrary hypergraphs, thereby establishing a general framework for recovering the hypergraph structure from symmetric simplicial set data.
$[v_{0}]_{e}$$[v_{1}]_{e}$$[v_{2}]_{e}$$[v_{3}]_{e}$$[v_{4}]_{e^{\prime}}$$[v_{4},v_{3}]_{e^{\prime}}$$[v_{2},v_{3}]_{e}$$[v_{2},v_{3}]_{e^{\prime}}$$[v_{1},v_{2},v_{0}]_{e}$
(a) $\Delta(H)$
$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$[[v_{0},v_{1},v_{2},v_{3}]_{e}]$$[[v_{2},v_{3},v_{4}]_{e^{\prime}}]$
(b) $H(\Delta(H))$
Figure 8: Illustrations of $\Delta(H)$, $H(\Delta(H))$ for a hypergraph $H=(\{v_{0},\cdots,v_{4}\},\{e,e^{\prime}\},f_{H})$ with $f_{H}(e)=\{v_{0},v_{1},v_{2},v_{3}\}$ and $f_{H}(e^{\prime})=\{v_{2},v_{3},v_{4}\}$: (a) $\Delta(H)$ is the collection of all possible oriented $n$-dimensional subrelations within $e,e^{\prime}$. Figure illustrates some 0, 1, and 2-simplices in $\Delta(H)$. (b) Figure illustrates a hypergraph $H(\Delta(H))$ consisting of five nodes and two nondegenerate maximal simplices serving as hyperedges. This reconstructed hypergraph $H(\Delta(H))$ is isomorphic to the original hypergraph $H$, demonstrating the information-preserving nature of $\Delta(H)$.
Definition 9
Let $H$ be a hypergraph. $H(\Delta(H))=(V,E,f)$ is the hypergraph with the following data:
  1. 1.
    $V:=V(H)$
  2. 2.
    $E$ := A collection of equivalence classes of maximal nondegenerate $n$-simplices for $n>0$.
  3. 3.
    For $[[v_{i_{0}},\cdots,v_{i_{\|e\|}}]_{e}]\in E$,
    $f([[v_{i_{0}},\cdots,v_{i_{\|e\|}}]_{e}]):=\{v_{i_{0}},\cdots,v_{i_{\|e\|}}\}.$
Proposition 1
For any hypergraph $H$, there exists a canonical isomorphism $H(\Delta(H))\cong H$ induced by the bijective correspondence between hyperedges and equivalence classes of maximal non-degenerate simplices.
Proof 1
Define a map $b:E\to E(H)$ by
$b([[v_{i_{0}},\cdots,v_{i_{\|e\|}}]_{e}]):=e.$
Suppose $e^{\prime}\in E(H)$ with $f_{H}(e^{\prime})=\{v_{0},\cdots,v_{\|e^{\prime}\|}\}$.
  • $[[v_{0},\cdots,v_{\|e^{\prime}\|}]_{e^{\prime}}]\in b^{-1}(e^{\prime})$ implies $b$ is surjective.
  • Suppose
    $b([[v_{i_{0}},\cdots,v_{i_{\|e^{\prime}\|}}]_{e^{\prime}}])=b([[v_{j_{0}},\cdots,v_{j_{\|e^{\prime}\|}}]_{e^{\prime}}])=e^{\prime}.$
    By definition of $b$, this implies
    $\{v_{i_{0}},\cdots,v_{i_{\|e^{\prime}\|}}\}=\{v_{j_{0}},\cdots,v_{j_{\|e^{\prime}\|}}\}$
    and $[[v_{i_{0}},\cdots,v_{i_{\|e^{\prime}\|}}]_{e^{\prime}}]=[[v_{j_{0}},\cdots,v_{j_{\|e^{\prime}\|}}]_{e^{\prime}}]$.
    Hence $b$ is injective.
Direct computations show that
$((\mathrm{Id}_{V(H)})_{*}\circ f)$ $([[v_{0},\cdots,v_{{\|e^{\prime}\|}}]_{e^{\prime}}])$
$=(f_{H}\circ b)([[v_{0},\cdots,v_{{\|e^{\prime}\|}}]_{e^{\prime}}])$
=$=\{v_{0},\cdots,v_{{\|e^{\prime}\|}}\}$
and
$(((\mathrm{Id}_{V(H)})^{-1})_{*}\circ f_{H})(e^{\prime})$ $=(f\circ b^{-1})(e^{\prime})$
=$=\{v_{0},\cdots,v_{\|e^{\prime}\|}\}.$
Therefore, $(\mathrm{Id}_{V(H)},b):H(\Delta(H))\to H$ is a hypergraph isomorphism. $\square$
This connection is particularly valuable, as $\Delta(H)$ possesses a richer structural framework compared to the original hypergraph $H$. The enhanced complexity of $\Delta(H)$ allows for a more nuanced analysis of the underlying hypergraph structure without losing the original hypergraph information $H$. In real-world hypergraphs, node features are often the only provided data. To leverage these features effectively, it is critical to explicitly compute the degree 0 sheaf Laplacian, which models how node features diffuse to adjacent nodes through the hypergraph structure.
Proposition 2
Suppose $H=(V(H),E(H),f_{H})$ is a hypergraph and $\mathcal{F}$ is a cellular sheaf on $\Delta(H)$. For a 0-cochain $\mathbf{x}$, the $[v]_{v}$-component of the degree 0 sheaf Laplacian is
$\sum_{\begin{subarray}{c}\{w\mid w,v\in f_{H}(e)\ \textrm{for}\ e\}\\ \{e\mid v,w\in f_{H}(e)\}\end{subarray}}\mathcal{F}^{*}([v]_{v}\unlhd[v,w]_{e})\mathcal{F}([v]_{v}\unlhd[v,w]_{e})(x_{v})$
$+$ $\sum_{\begin{subarray}{c}\{w\mid w,v\in f_{H}(e)\ \textrm{for}\ e\}\\ \{e\mid v,w\in f_{H}(e)\}\end{subarray}}\mathcal{F}^{*}([v]_{v}\unlhd[w,v]_{e})\mathcal{F}([v]_{v}\unlhd[w,v]_{e})(x_{v})$
-- $\sum_{\begin{subarray}{c}\{w\mid w,v\in f_{H}(e)\ \textrm{for}\ e\}\\ \{e\mid v,w\in f_{H}(e)\}\end{subarray}}\mathcal{F}^{*}([v]_{v}\unlhd[v,w]_{e})\mathcal{F}([w]_{w}\unlhd[v,w]_{e})(x_{w})$
-- $\sum_{\begin{subarray}{c}\{w\mid w,v\in f_{H}(e)\ \textrm{for}\ e\}\\ \{e\mid v,w\in f_{H}(e)\}\end{subarray}}\mathcal{F}^{*}([v]_{v}\unlhd[w,v]_{e})\mathcal{F}([w]_{w}\unlhd[w,v]_{e})(x_{w}).$
Proof 2
  • $[v]_{v},[w]_{w}$ are upper adjacent whose cofacets are $[v,w]_{e},[w,v]_{e}$ for any $e\in E(H)$ such that $v,w\in f_{H}(e)$. Moreover,
    1. 1.
      $[[v]_{v}:[v,w]_{e}]=[[w]_{w}:[w,v]_{e}]=-1$
    2. 2.
      $[[w]_{w}:[v,w]_{e}]=[[v]_{v}:[w,v]_{e}]=+1$.
  • $[v]_{v},[v]_{v}$ are upper adjacent whose cofacets are $[v,w]_{e},[w,v]_{e}$ for any $e\in E(H)$ such that $v,w\in f_{H}(e)$. Moreover,
    1. 1.
      $[(v)_{V}:(v,w)_{V}]=-1$
    2. 2.
      $[(v)_{V}:(w,v)_{V}]=+1$.
Hence the degree 0 sheaf Laplacian easily follows by Equation (8). $\square$
Given a graph $G$, we have two normalized sheaf Laplacians : Equation (1) and Proposition 2. The next proposition says twos are closely related.
Proposition 3
Suppose $G$ is a graph and $\mathcal{F}$ is a cellular sheaf on $G$. Then $\mathcal{F}$ induces a cellular sheaf $\widehat{\mathcal{F}}$ on $\Delta(G)$ satisfying
^$\mathcal{L}^{0}_{\widehat{\mathcal{F}}}=\mathcal{L}_{\mathcal{F}}.$
Proof 3
Define a cellular sheaf $\widehat{\mathcal{F}}$ on $\Delta(G)$ as follows. First, stalks are defined by
^$\widehat{\mathcal{F}}([v_{i_{0}},\cdots,v_{i_{n}}]_{e}):=\begin{cases}\mathcal{F}(v)\textrm{\ if $\{v_{i_{0}},\cdots,v_{i_{n}}\}=\{v\}$ }\\ \mathcal{F}(e)\textrm{\ otherwise.}\end{cases}$
Linear map ^$\widehat{\mathcal{F}}([v_{i_{0}},\cdots,\widehat{v_{i_{l}}},\cdots,v_{i_{n}}]_{e}\unlhd[v_{i_{0}},\cdots,v_{i_{n}}]_{e})$ is defined by
  • $\mathcal{F}(v\unlhd e)$ if $\{v_{i_{0}},\cdots,\widehat{v_{i_{l}}},\cdots,v_{i_{n}}\}\neq\{v_{i_{0}},\cdots,v_{i_{n}}\}$
  • $\mathrm{Id}$ otherwise.
Note that
$C^{0}(\Delta(G),\widehat{\mathcal{F}})=C^{0}(G,\mathcal{F})=\underset{{v\in V(G)}}{\oplus}\mathcal{F}(v).$
Hence $L^{0}_{\widehat{\mathcal{F}}},D^{0}_{\widehat{\mathcal{F}}},L_{\mathcal{F}},D_{\mathcal{F}}$ are linear maps from $C^{0}(G,\mathcal{F})$ to $C^{0}(G,\mathcal{F})$. When $v\unlhd e$, there are two cofacets $[v,w]_{e},[w,v]_{e}$ of $[v]_{v}$. Equations (1), (3),(9) and Proposition 2 imply
$L^{0}_{\widehat{\mathcal{F}}}=2L_{\mathcal{F}},D^{0}_{\widehat{\mathcal{F}}}=2D_{\mathcal{F}}.$
Therefore,
^$\mathcal{L}^{0}_{\widehat{\mathcal{F}}}$ $=(D^{0}_{\widehat{\mathcal{F}}})^{-\frac{1}{2}}L^{0}_{\widehat{\mathcal{F}}}(D^{0}_{\widehat{\mathcal{F}}})^{-\frac{1}{2}}$
$=(2D_{{\mathcal{F}}})^{-\frac{1}{2}}(2L_{{\mathcal{F}}})(2D_{{\mathcal{F}}})^{-\frac{1}{2}}$
$=(D_{{\mathcal{F}}})^{-\frac{1}{2}}(L_{{\mathcal{F}}})(D_{{\mathcal{F}}})^{-\frac{1}{2}}$
$=\mathcal{L}_{\mathcal{F}}$
and it finishes the proof. $\square$
Proposition 3 implies the normalized degree $k$ sheaf Laplacian on $\Delta(G)$ is a generalization of the normalized sheaf Laplacian on $G$.
IV-C Neural sheaf diffusion on hypergraph
In this subsection, we develop a degree $k$ NSD model on a symmetric simplicial set via the normalized degree $k$ sheaf Laplacian. By instantiating this framework on a symmetric simplicial set induced from hypergraph, we derive HNSD, enabling structure-aware learning on higher-order networks.
Let $X$ be a symmetric simplicial set, $\mathcal{F}$ be a cellular sheaf on $X$ of degree $k+1$ such that $\mathcal{F}(\sigma)=\mathbb{R}^{d}$ for any simplex $\sigma$. Let $\mathbf{x}_{1},\cdots,\mathbf{x}_{f}$ be $k$-cochains and $\mathbf{X}\in\mathbb{R}^{|X_{k}|d\times f}$ be a matrix whose $i$th column is $\mathbf{x}_{i}$ for $1\leq i\leq f$. Degree $k$ sheaf diffusion of $\mathbf{X}$ on $(X,\mathcal{F})$ is time dependent matrices $\mathbf{X}(t)$ describing the diffusion of $\mathbf{X}$ along $\mathcal{F}$ governed by
$\mathbf{X}(0)=\mathbf{X},\ \dot{\mathbf{X}}(t)=-\mathcal{L}^{k}_{\mathcal{F}}(\mathbf{X}(t)).$ (12)
Equation (12) is discretized via the explicit Euler scheme with unit step-size
$\mathbf{X}_{t+1}=\mathbf{X}_{t}-\mathcal{L}^{k}_{\mathcal{F}}(\mathbf{X}_{t}).$
A general layer of degree $k$ NSD on $X$ is defined by
$\mathbf{X}_{t+1}=\mathbf{X}_{t}-\alpha\left(\mathcal{L}^{k}_{\mathcal{F}(t)}(\mathbf{Id}\otimes\mathbf{W}^{t}_{1})\mathbf{X}_{t}\mathbf{W}^{t}_{2}\right)$ (13)
where
  • $t\in\mathbb{N}$ is layer
  • $\alpha$ is nonlinear function
  • $\mathbf{W}^{1}_{t}\in\mathbb{R}^{d\times d},\mathbf{W}^{2}_{t}\in\mathbb{R}^{f_{t}\times f_{t+1}}$ are learnable matrices at layer $t$
  • $\mathcal{F}(t)$ is cellular sheaf on $X$ at layer $t$.
The degree $k$ NSD on symmetric simplicial sets generalizes NSD on graphs [17] through Proposition 3. For a hypergraph $H$ with induced symmetric simplicial set $\Delta(H)$, we learn the cellular sheaf $\mathcal{F}$ on $\Delta(H)$ of degree 1 via
$\mathcal{F}([v]_{v}\unlhd[v,w]_{e})=\mathrm{MLP}(x_{v}\parallel x_{[v,w]_{e}})$ (14)
where
$x_{[v,w]_{e}}=\sigma^{\prime}\left(M\sigma\left(W^{T}\left[\begin{array}[]{c}x_{v}\\ 1\end{array}\right]\odot\left[\begin{array}[]{c}x_{w}\\ 1\end{array}\right]\right)\right)$
with $\sigma=\mathrm{ReLU}$, $\sigma^{\prime}=\tanh$ [38].
The HNSD architecture on $H$ is defined as the degree 0 NSD on $\Delta(H)$ with sheaf learning governed by Equation (14). This formulation preserves the theoretical guarantees of NSD while enabling structural learning on higher-order relations through the symmetric simplicial set framework.
V Numerical experiments
In this section, we present the experimental results and provide a comprehensive analysis of the performance achieved by the proposed method. To thoroughly assess its effectiveness, evaluations were conducted across five diverse real-world datasets.
V-A Experiment settings
V-A1 Dataset
In our experiments, we utilized five real-world datasets. The Cora and Citeseer datasets are citation networks, while the Cora CA and DBLP CA datasets are co-authorship networks [29]. Citation datasets capture relationships between scientific publications based on citation patterns, where hypervertices represent individual papers and hyperedges link groups of papers that are co-cited by another publication. Similarly, the co-authorship datasets reflect relationships between papers based on shared authorship. Here, hypervertices represent individual papers, and hyperedges connect papers authored by the same researcher(s). Additionally, we used the Senate datasets [39], which represent legislative cosponsorship networks within the United States Congress. These datasets model the relationships between legislators and the bills they sponsor or cosponsor. In this case, hypervertices represent individual senators, while hyperedges denote legislative bills that link sponsors to their cosponsors. Each hypervertex contains attributes indicating the legislator’s political party affiliation. Detailed characteristics of each dataset are summarized in Table II.
TABLE II: Dataset statistics. CE homophily: homophily score [40] from clique expansion. Avg. HE size: average hyperedge size.
Cora Citeseer Cora-CA DBLP-CA Senate
# hypernodes 2708 3312 2708 41302 282
# hyperedges 1579 1079 1072 22363 315
# classes 7 6 7 6 2
Avg. HE size 3.03 3.20 4.28 4.45 17.17
CE Homophily 0.897 0.893 0.803 0.869 0.498
V-A2 Task and evaluation
TABLE III: Node classification accuracy on five benchmark hypergraph datasets
Cora Citeseer Cora-CA DBLP-CA Senate AVG.
CEGCN 75.32 ± 1.69 71.43 ± 1.34 76.68 ± 1.30 87.19 ± 0.30 48.17 ± 3.68 71.74
HNHN 76.36 ± 1.92 72.64 ± 1.57 77.19 ± 1.49 86.78 ± 0.29 50.85 ± 3.35 72.76
LEGCN 72.23 ± 1.60 71.84 ± 1.17 72.23 ± 1.60 84.26 ± 0.40 73.24 ± 10.29 74.76
HyperGCN 74.19 ± 1.41 69.42 ± 3.49 70.00 ± 3.74 86.78 ± 2.39 53.66 ± 6.35 70.78
AllDeepSets 76.88 ± 1.80 70.83 ± 1.63 81.97 ± 1.50 91.27 ± 0.27 48.17 ± 5.67 73.82
SheafHyperGNN 77.80 ± 2.24 73.93 ± 1.06 81.65 ± 1.50 88.93 ± 0.66 74.65 ± 5.90 79.39
HNSD(OURS) 78.02 ± 1.78 74.02 ± 1.74 82.64 ± 1.17 89.67 ± 0.40 78.45 ± 5.87 80.56
Our experiment focuses on node classification in hypergraphs. Each node in the hypergraph is assigned a label corresponding to one of several classes, depending on the characteristics of the dataset. The goal of our task is to predict the labels of unknown nodes based on the given hypergraph structure. To achieve this, we partition each dataset into training, validation, and test sets with a ratio of 0.5:0.25:0.25. The neural network is first trained using the training dataset. The validation dataset is used to monitor the model’s performance, and the final test accuracy is measured at the epoch where the validation error is minimized. The evaluation metric used in our experiment is accuracy, which represents the proportion of correctly classified nodes among all tested nodes. To ensure the reliability of our results, each experiment is conducted 10 times, and we report the average accuracy across these runs.
V-A3 Baseline
In our evaluation, we compared our approach against several established baseline methods. CEGCN [5] integrates the Clique Expansion (CE) technique with Graph Convolutional Networks (GCNs). It transforms hypergraphs into standard graphs by converting each hyperedge into a fully connected clique, upon which conventional GCNs are applied. HNHN [28] is a hypergraph neural network that alternates between updating hyperedge and hypervertex representations using a normalization scheme, thereby effectively capturing higher-order dependencies inherent in hypergraph structures. LEGCN [5] incorporates label embeddings into the graph convolutional framework by transforming hypergraphs into standard graphs where node features are augmented with label information, enhancing the representation of relational structures. HyperGCN [29] extends GCNs to the hypergraph domain by introducing a hypergraph Laplacian and reducing hyperedges into weighted pairwise edges in a selective manner, enabling the model to retain essential high-order relationships. AllDeepSets [41] models hypergraph message passing through two permutation-invariant multiset functions, both instantiated via Deep Sets. These functions, implemented as multilayer perceptrons (MLPs) with summation-based aggregation, provide universal approximation capabilities for multiset functions. SheafHyperGNN [32] builds upon sheaf theory to enrich GNNs by incorporating both local and global contexts through a hypergraph-based framework. It introduces a sheaf-theoretic linear diffusion mechanism to effectively model higher-order interactions within hypergraphs.
V-B Results and Analysis
V-B1 Overall Performance Comparison
Table III presents the classification accuracy of the proposed HNSD model across five benchmark datasets—Cora, Citeseer, Cora-CA, DBLP-CA, and Senate—spanning citation, co-authorship, and legislative cosponsorship networks. HNSD is evaluated alongside several competitive baselines, with accuracy serving as the primary metric for assessing node classification performance.
On citation networks (Cora and Citeseer), HNSD achieves marginal yet consistent improvements of 0.22% and 0.09%, respectively, over the best-performing baselines. These results indicate that HNSD effectively captures fine-grained dependencies among co-cited nodes by leveraging its sheaf-based diffusion mechanism.
More substantial gains are observed in co-authorship networks. On the Cora-CA dataset, HNSD surpasses AllDeepSets by 0.67%, while also outperforming all baselines on DBLP-CA. This suggests that the node-sensitive design of HNSD is particularly well-suited to modeling collaborative structures inherent in co-authorship hypergraphs.
The most pronounced improvement is reported on the Senate dataset, where HNSD exceeds SheafHyperGNN by 3.80%. Given the heterogeneity of node attributes and the sparsity of connections in this dataset, this result highlights the model’s robustness in challenging, real-world scenarios.
Overall, HNSD achieves the highest average accuracy of 80.56%, outperforming SheafHyperGNN and AllDeepSets by margins of 1.17% and 6.74%, respectively. These consistent improvements across diverse domains validate the effectiveness of incorporating sheaf-theoretic structures with node-adaptive diffusion.
V-B2 Comparison with SheafHyperGNN
A key innovation of HNSD lies in its rigorous mathematical grounding. While SheafHyperGNN introduces sheaf-based diffusion for hypergraph learning, our method advances this approach by constructing a sheaf Laplacian on hypergraphs formalized as symmetric simplicial sets. This abstraction offers a principled framework that more faithfully captures higher-order relationships.
This enhanced theoretical foundation directly translates into empirical gains. Across all five datasets, HNSD consistently outperforms SheafHyperGNN, including a significant margin of 3.80% on the Senate dataset. These results underscore the practical benefits of our design, demonstrating that a formally structured sheaf-based framework yields measurable performance advantages.
V-B3 Strength in Heterophilic Hypergraphs
The Senate dataset is a prime example of a heterophilic hypergraph, where nodes connected by hyperedges typically exhibit highly dissimilar attributes. This low CE homophily value indicates that, unlike traditional graphs where connected nodes tend to share similar features, the relationships between connected nodes in the Senate dataset are more complex and varied. The dataset contains political entities with diverse attributes (e.g., political affiliations, and voting behaviors), and the connections between them are sparse, making it challenging for conventional models that assume homophily.
In this context, HNSD excels by leveraging a sheaf-based diffusion mechanism that is particularly well-suited for heterophilic settings. Unlike traditional methods that rely on fixed node features, HNSD dynamically learns the sheaf stalks through an MLP layer, enabling the model to incorporate high-order structural information into the node representations. This allows the model to capture intricate relationships between nodes and hyperedges, even when the nodes themselves are dissimilar in attributes. The restriction map within the sheaf framework further enhances information propagation, ensuring that even dissimilar nodes within the same hyperedge can influence each other effectively.
Thus, HNSD’s ability to learn node-sensitive dynamics and model higher-order relationships provides a significant advantage in the Senate dataset. By integrating sheaf-theoretic principles with node-specific diffusion, HNSD outperforms existing methods in handling heterophilic interactions, leading to a 3.80% performance boost over the second-best model, SheafHyperGNN. This demonstrates that HNSD is particularly adept at handling real-world hypergraphs with irregular topologies and heterogeneous attributes, making it a strong contender for complex datasets like the Senate dataset.
VI Conclusion
In this work, we propose a new hypergraph neural network HNSD, a principled generalization of NSD from graphs to hypergraphs. Our framework constructs symmetric simplicial sets directly from hypergraphs, encoding all oriented subrelations within each hyperedge as ordered tuples. This construction preserves the original hypergraph’s structural information. By canonically defining adjacency via shared facets, our normalized degree 0 sheaf Laplacian captures non-trivial interactions between nodes, avoiding the sparsity-induced degeneracy of traditional degree 0 hypergraph Laplacian. We rigorously prove that normalized degree 0 sheaf Laplacian reduces exactly to the traditional normalized sheaf Laplacian on graphs, establishing our framework as the natural generalization of sheaf-based learning to higher-order domains. Empirically, HNSD achieves competitive performance across established hypergraph benchmarks, validating its efficacy as the first structure-preserving sheaf-theoretic extension to hypergraph neural networks. An interesting direction for future research is the extension of HNSD to directed hypergraphs, allowing for richer and more nuanced relational modeling.